# Planet-Planet Scattering and ZLK Migration – The Dynamical History of HAT-P-11

### v1

@@ 3.3 Radius Evolution
It turns out that story we just elucidated is incomplete, since simple ZLK migration alone cannot explain HAT-P-11 b’s orbit. This is due to how relatively close-in HAT-P-11 c is. From Equation (7), the precession rate associated with the ZLK effect scales very strongly with $a_{\mathrm{c}}$. Since $a_{\mathrm{c}}$ is small, the precession associated with ZLK oscillations is strong, and thus it requires very high eccentricities for tides to break the ZLK cycles (e.g., in comparison to GJ 436). Inspection of Figure 5 provides a clearer picture. The contour that HAT-P-11 b presently lies on does not intersect the solid golden line in the phase space shown. While it does intersect the dark red line denoting suppression of ZLK oscillations due to GR precession, this mechanism would require suppression of ZLK oscillations at an eccentricity of just $\approx$0.5 at a semimajor axis where tidal damping is slow, and is not something that we see in our simulations. In order for HAT-P-11 b to have a constant $L$ path from tidal quenching of ZLK oscillations to its current orbit, we would have to extend Figure 5 well to the right, at which point HAT-P-11 b would participate in the scattering events discussed in Section 2. In other words, the quenching of ZLK by tidal effects in a reasonable initial parameter space cannot bring HAT-P-11 b to its present-day orbit via tidal circularization. We see that more realistic contours of constant $L$ terminate significantly closer to the star, sometimes by up to an order of magnitude. This tendency to overshoot HAT-P-11 b’s present-day orbit is insensitive to initial conditions and is confirmed with N-body simulations. Thus, the present-day configuration cannot be explained with traditional ZLK migration with fixed planet radius and equilibrium tides alone.
@@
This effect may play a significant role in the dynamics of the system, as from Equation (11) tidal precession scales as $\propto R_{\rm b}^{5}$. Indeed, we see from the dotted golden line in Figure 5 that increasing the radius of HAT-P-11 b by a factor of two breaks the ZLK oscillations significantly earlier, and now there does exist a reasonable contour of constant $L_{\rm b}$ that can deliver HAT-P-11 b to its present-day orbit. Coupled dynamical and thermal evolution of hot Jupiters were explored in the context of planet-planet scattering by Rozner et al. (2022); Glanz et al. (2022), who found that accounting for radius inflation significantly accelerated the formation and destruction of hot Jupiters and enhanced the production of warm Jupiters. Fully self-consistent coupled dynamical and thermal evolution in the context of ZLK migration has not been explored in detail. Given the uncertainties in our system, we employ a simplistic model for the radius inflation of HAT-P-11 b in this work, and we defer a more nuanced prescription for self-consistent thermal and orbital evolution to future work.
Millholland et al. (2020) generated 10,000 MESA models, which modify and build upon the publicly available MESA models for sub-Saturns developed by Chen & Rogers (2016). These models spanned four principle parameters: mass $M$, envelope mass fraction $f_{\mathrm{env}}$, flux from the host star $F$, and internal luminosity $\mathcal{L}$ [^1] Note that Millholland et al. (2020) parameterized the tidal luminosity with the factor $\Gamma\equiv\log_{10}\left[\frac{Q^{\prime}(1+\cos^{2}\epsilon)}{\sin^{2}\epsilon}\right]$, where $Q^{\prime}\equiv 3Q/2k_{2}$ is the reduced tidal quality factor and $\epsilon$ is the planetary obliquity. This is because both Millholland et al. (2020) and their progenitor study Millholland (2019) focused on obliquity tides. However, we are agnostic to the specific source of tides in this study, and as such require only the total luminosity deposited at the core-envelope interface.. Each model was then evolved forward in time for $10$ Gyr, after which a final radius $R$ value was reported. We fit a subset of the Millholland et al. (2020) models appropriate to HAT-P-11 b in order to construct a luminosity-radius relationship using a fourth-order polylogarithmic function:
@@
where $A=5.23\times 10^{-4}$, $B=3.37\times 10^{-2}$, $C=0.801$, $D=8.357$, and $E=3.596\times 10^{1}$. Here the orbit-averaged tidal luminosity of the planet is given (Mardling & Lin, 2002):
@@
We now use our fit to the MESA models to evolve HAT-P-11 b’s radius during the simulations. In the interest of computation time and as our expression is orbit-averaged, we do not recompute the planet’s radius every timestep. Rather, radius evolution is computed once per orbit of HAT-P-11 b. Once per orbit, the orbit-averaged luminosity is calculated and fed into our luminosity-radius relation Eq. (12), and the radius of HAT-P-11 b is correspondingly updated. We again emphasize that our prescription for the thermal radius evolution of the planet is a first-order approximation, where we assume the planet’s radius responds effectively instantaneously to the tidal luminosity. A more realistic prescription would include a time-delayed response to the tidal heating, and a fully self-consistent numerical prescription could involve codes that couple the dynamical evolution of the system to the planet’s thermal evolution (e.g. see Glanz et al. (2022), who coupled outputs from REBOUND and MESA). While we do not use these nuanced prescriptions in the present work, we argue that our first-order approximation qualitatively achieves similar results. As our orbital code tracks tidal luminosity self-consistently, our planet will ultimately be inflated to the same degree, and the ZLK cycles will be broken at similar times. The improvement which would likely produce the largest discrepancy between our model is our choice of equilibrium tides, but we defer treatment of the more realistic dynamical tide model to future work.
@@ 3.4 N-body simulations
We investigate the present-day orbit of HAT-P-11 b via ZLK migration through a suite of N-body simulations using REBOUND. The setup of our numerical simulations are as follows. We initialize HAT-P-11A and c in their present-day orbital configurations. We select the best-fit values for the mass, semimajor axis and eccentricity of HAT-P-11 c from Table 1. The stellar obliquity of HAT-P-11 c is less constrained; we draw from a uniform distribution $\psi_{\mathrm{c,i}}\in\{$$,$$\}$ 36.17  ° , 65.24  ° }, where the lower bound is selected from the $3\sigma$ confidence interval from the observational posteriors, and the upper bound from the $2\sigma$ confidence interval of our four-body scattering simulations. HAT-P-11 b is initialized with orbital elements randomly drawn from uniform distributions informed from the $1\sigma$ distributions obtained in our four-body scattering simulations: $a_{\mathrm{b}}=\{0.27\ \mathrm{AU},0.52\ \mathrm{AU}\}$, $e_{\mathrm{b}}=\{0.04,0.627\}$ and $\psi_{\mathrm{b}}=\{12.9^{\circ},64.7^{\circ}\}$. We initialize the mass with the best-fit value from Table 1, and the radius with Equation (12).
@@
We also model the effects of general relativity and tides. We use REBOUNDx (Tamayo et al., 2020) to model both. For GR precession, we use the gr prescription (Anderson et al., 1975), appropriate for systems with a dominant central mass. We consider self-consistent spin, tidal and orbital evolution using the tides_spin effect (Lu et al., 2023), which provides the ability to evolve the evolution of the stellar spin axis for consistent stellar obliquity tracking as well. For planet b, we use fiducial parameters of $k_{2}=0.5$, $I=0.25$ appropriate for a Uranus-like planet (Yoder, 1995; Lainey, 2016), and we initialize its spin state as both pseudo-synchronized and aligned with the orbit. Planet c is considered a point particle, and the effects of structure are not accounted for.
@@
Tides on the host star and planet c are not considered, but we set the value of planet b’s tidal parameter to $\tau_{\rm b}=10^{-4}$ in the default REBOUND units, where $G=1$. This is an unrealistically high value of $\tau$, chosen to minimize computation time: direct N-body simulations with a realistic $\tau$ value proved to be computationally infeasible. This is a common practice in numerical simulations (e.g Bolmont et al., 2015; Becker et al., 2020). Scaling $\tau$ does not change the qualitative behavior of the system. Neither the ZLK timescale nor the tidal precession rate, Equations (5) and (11) respectively, depend on the value of the tidal parameter. The only dependence on $\tau$ is the rate of orbital energy drained from tides. Thus, our scaled-up value of $\tau$ serves only to drain more energy from the planet’s orbit every ZLK cycle (therefore decreasing the semimajor axis more) and increase the rate at which HAT-P-11 b circularizes once the ZLK oscillations are ultimately quenched. These both only serve to linearly modify the timescale in which the planet reaches the final stage of evolution.
@@
We now briefly discuss what realistic values of $\tau$ may be. While the values of tidal parameters in exoplanets are very poorly constrained, estimates of Neptune’s $\tau\sim 10^{-7}$ are based on the orbit of its satellites Proteus and Larissa (Zhang & Hamilton, 2007). Therefore, the tidal timescale in the N-body simulations is a factor $\sim$103 longer than the simulation timescale.
@@ 3.5 Simulation Results
Figure 7: Results from a population synthesis of ZLK simulations. We consider $68$ simulations with radius evolution (orange) and $64$ with fixed radius (blue). We compare distributions of semimajor axis, stellar obliquity and mutual inclination. We see that the fixed radius model cannot generate HAT-P-11 b’s present-day semi-major axis, but the model with tidally-driven radius inflation can.
@@
Figure 7 shows distributions of stellar obliquity and eccentricity for a population synthesis of $100$ systems. We initialize each simulation as described in the previous section and run until the ZLK cycles are quenched and the orbit has circularized to the present-day eccentricity $e_{\mathrm{b}}=0.218$. At this point, the simulation is halted. To illustrate the effects of radius inflation, we also run the same initial condition but fix the radius of HAT-P-11 b to its present-day value throughout the course of the simulation. We allow each simulation to run for 10 Myr, which corresponds to 10 Gyr in when time is rescaled appropriately. $68$ simulations with radius inflation are able to quench ZLK oscillations and circularize in this timeframe, with the remaining $32$ either never initiating ZLK oscillations or being unable to quench them in time. For the simulations with fixed radius, $64$ simulations quench and circularize in time. For the population synthesis that considers radius inflation, we see that while the best-fit observational values are not the most favored there is nonetheless a high proportion of simulations that match the constraints, and we do not require fine-tuning to reproduce our posteriors. The population synthesis without radius inflation does not exhibit significant differences in the stellar obliquity or mutual inclination distributions. However, as expected this model greatly underpredicts the value the semimajor axis of HAT-P-11 b: all simulations predict an orbit with $a_{\mathrm{b}}<0.045$ AU (consistent with the implication of Figure 5).
@@
Figure 8: Fiducial ZLK migration of HAT-P-11 b. From top to bottom, HAT-P-11 b’s semimajor axis, eccentricity, stellar obliquity and radius are shown with the colored lines. The black dots represent HAT-P-11 b’s measured orbital parameters, with the error bar indicating the $1\sigma$ confidence interval. We posit that HAT-P-11 b is still in the tidal circularization stage after the suppression of ZLK oscillations. There is excellent agreement between our simulation results and observed values.
@@
Figure 8 shows one of our simulations in detail. Table 2 lists the initial orbital parameters of the simulation, as well as the final values of of interest. During the first phase of the simulation, ZLK oscillations are active and the planet experiences the characteristic coupled eccentricity and inclination oscillations. At each high-eccentricity epoch, tides act on the planet to lower its semimajor axis. The radius evolution of the planet is also plotted, and we see that during each high-eccentricity epoch the planet inflates to over twice its present-day radius. Note that the minimum pericenter distance (${\sim}0.02\ \mathrm{AU}$) is well outside twice of the tidal disruption radius ($r_{t}=R_{p}(M_{*}/M_{p})^{1/3}{\sim}0.008\ \mathrm{AU}$), so we do not include mass loss due to tidal stripping in our model, as the effects are weak (Guillochon et al., 2011). Later, the ZLK oscillations are damped during a high eccentricity epoch. Tides continue to act to circularize the planet’s orbit. In the meantime, the orbit normal of HAT-P-11 b precesses about the invariant plane of the system (essentially HAT-P-11 c’s orbit normal), as expected, resulting in periodic stellar obliquity oscillations. As expected, the stellar spin axis and the orbit of HAT-P-11 c do not change significantly over the course of the integration.
@@
We comment again on the interpretation of timescales, given our choice of $\tau$ in order to speed up computations. Please refer to Section 3.4 for justification regarding the choice of an enhanced dissipation rate. The enhanced dissipation rate affects the phase when tidal dissipation dominates, and thus we scale the timescales on the right-hand side of the panel to the realistic values accordingly. Note that with a more realistic tidal time-lag, we would expect to see more stellar obliquity cycles on the right-hand side of the plot when tides dominate (since the orbital precession frequency does not depend on tidal dissipation and shouldn’t be scaled). In addition, we would expect to have more ZLK cycles on the left-hand side, since each cycle leads to a weaker semi-major axis decay. However, the qualitative behavior of the evolution is the same.
@@
The age of the HAT-P-11 system is not well-constrained: Yee et al. (2018) propose an age of $6.5^{+5.9}_{-4.1}$ Gyr, while Morton et al. (2016) present an estimate of $2.69^{+2.88}_{-1.24}$ Gyr. Our results are consistent with the system age. Note that the timescale of ZLK migration varied by around an order of magnitude, depending on initial conditions.
@@
$\mathrm{R}$ ($\mathrm{R_{Earth}}$) 3.60 4.38 N/A N/A
$a$ (AU) 0.496 0.0525 4.192 4.194
$e$ 0.31 0.21 0.56 0.56
$\psi$ $(^{\circ})$ 49.0 133.2 55.6 55.0
Note. — Relevant input parameters and output values of the REBOUND simulation shown in Figure 8. The final values are the values at the time corresponding to the black dot in Figure 8, which is where $a_{\mathrm{b}}$ is equal to the best fit value. The orbit of HAT-P-11 c does not change significantly over the course of the simulation.
The present-day observational constraints on HAT-P-11 b’s orbit are also shown on Figure 8. Our simulation is able to match all values very well. We propose that HAT-P-11 b is currently in this final stage of evolution and is still in the process of tidal circularization. We conclude that ZLK migration coupled with thermally driven radius evolution is capable of explaining the present-day orbit of HAT-P-11 b remarkably well. Again, we emphasize that we did not perform an exhaustive parameter space analysis of this system. Rather, our aim is to qualitatively show that such a process is capable of reproducing HAT-P-11 b’s orbit.

### v2

@@ 3.3 Radius Evolution
It turns out that the story we just elucidated is incomplete, since simple ZLK migration alone cannot explain HAT-P-11 b’s orbit. This is due to how relatively close-in HAT-P-11 c is. From Equation (8), the precession rate associated with the ZLK effect scales very strongly with $a_{\mathrm{c}}$. Since $a_{\mathrm{c}}$ is small, the precession associated with ZLK oscillations is strong, and thus very high eccentricities are required for tides to break the ZLK cycles (e.g., in comparison to GJ 436). Inspection of Figure 5 provides a clearer picture. The contour that HAT-P-11 b presently lies on does not intersect the solid golden line in the phase space shown. While it does intersect the dark red line denoting suppression of ZLK oscillations due to GR precession, this mechanism would require suppression of ZLK oscillations at an eccentricity of just $\approx$0.5 at a semimajor axis where tidal damping is slow, and is not something that we see in our simulations. In order for HAT-P-11 b to have a constant $L$ path from tidal quenching of ZLK oscillations to its current orbit, we would have to extend Figure 5 well to the right. We find that initializing HAT-P-11 b at these larger orbital separations typically results in dynamical instability, and often ejection of HAT-P-11 b. In other words, the quenching of ZLK by tidal effects in a reasonable initial parameter space cannot bring HAT-P-11 b to its present-day orbit via tidal circularization. We see that more realistic contours of constant $L$ terminate significantly closer to the star, sometimes by up to an order of magnitude. This tendency to overshoot HAT-P-11 b’s present-day orbit is insensitive to initial conditions and is confirmed with N-body simulations. Thus, the present-day configuration cannot be explained with traditional ZLK migration with fixed planet radius and equilibrium tides alone.
@@
This effect may play a significant role in the dynamics of the system, as from Equation (12) tidal precession scales as $\propto R_{\rm b}^{5}$. Indeed, we see from the dotted golden line in Figure 5 that increasing the radius of HAT-P-11 b by a factor of two breaks the ZLK oscillations significantly earlier, and now there does exist a reasonable contour of constant $L_{\rm b}$ that can deliver HAT-P-11 b to its present-day orbit. Coupled dynamical and thermal evolution of hot Jupiters were explored in the context of planet-planet scattering by Rozner et al. (2022); Glanz et al. (2022), who found that accounting for radius inflation significantly accelerated the formation and destruction of hot Jupiters and enhanced the production of warm Jupiters. Petrovich (2015) investigated ZLK migration for initially inflated planets which slowly shrink down. However, fully self-consistent coupled dynamical and thermal evolution in the context of ZLK migration has not been explored in detail. Given the uncertainties in our system, we employ a simplistic model for the radius inflation of HAT-P-11 b in this work, and we defer a more nuanced prescription for self-consistent thermal and orbital evolution to future work.
Millholland et al. (2020) generated 10,000 MESA models (Paxton et al., 2011), which modify and build upon the publicly available MESA models for sub-Saturns developed by Chen & Rogers (2016). These models spanned four principal parameters: mass $M$, envelope mass fraction $f_{\mathrm{env}}$, flux from the host star $F$, and internal luminosity $\mathcal{L}$ [^2] Note that Millholland et al. (2020) parameterized the tidal luminosity with the factor $\Gamma\equiv\log_{10}\left[\frac{Q^{\prime}(1+\cos^{2}\epsilon)}{\sin^{2}\epsilon}\right]$, where $Q^{\prime}\equiv 3Q/2k_{2}$ is the reduced tidal quality factor and $\epsilon$ is the planetary obliquity. This is because both Millholland et al. (2020) and their progenitor study Millholland (2019) focused on obliquity tides. However, we are agnostic to the specific source of tides in this study, and as such require only the total luminosity deposited at the core-envelope interface.. Each model was then evolved forward in time for $10$ Gyr, after which a final radius $R$ was reported. We fit a subset of the Millholland et al. (2020) models appropriate to HAT-P-11 b in order to construct a luminosity-radius relationship using a fourth-order polylogarithmic function:
@@
where $A=5.23\times 10^{-4}$, $B=3.37\times 10^{-2}$, $C=0.801$, $D=8.357$, and $E=3.596\times 10^{1}$. Here the orbit-averaged tidal luminosity of the planet is given by (Mardling & Lin, 2002):
@@
We now use our fit to the MESA models to evolve HAT-P-11 b’s radius during the simulations. In the interest of computation time and as our expression is orbit-averaged, we do not recompute the planet’s radius every timestep. Rather, radius evolution is computed once per orbit of HAT-P-11 b. Once per orbit, the orbit-averaged luminosity is calculated and fed into our luminosity-radius relation Eq. (13), and the radius of HAT-P-11 b is correspondingly updated. We again emphasize that our prescription for the thermal radius evolution of the planet is a first-order approximation, where we assume the planet’s radius responds effectively instantaneously to the tidal luminosity. A more realistic prescription would include a time-delayed response to the tidal heating, and a fully self-consistent numerical prescription could involve codes that couple the dynamical evolution of the system to the planet’s thermal evolution. For example, see Glanz et al. (2022) who coupled outputs from MESA (Paxton et al., 2011) and AMUSE (Portegies Zwart et al., 2009). While we do not use these nuanced prescriptions in the present work, we assert that our first-order approximation qualitatively achieves similar results. As our orbital code tracks tidal luminosity self-consistently, our planet will ultimately be inflated to the same degree, and the ZLK cycles will be broken at similar times. The improvement which would likely produce the largest discrepancy between our model is our choice of equilibrium tides, but we defer treatment of the more realistic dynamical tide model to future work.
@@ 3.4 N-body simulations
We investigate the present-day orbit of HAT-P-11 b via ZLK migration through a suite of N-body simulations using REBOUND. The setup of our numerical simulations are as follows. We initialize HAT-P-11 A and c in their present-day orbital configurations. We select the best-fit values for the mass, semimajor axis and eccentricity of HAT-P-11 c from Table 1. The spin-orbit misalignment of HAT-P-11 c is less constrained; we draw from a uniform distribution $\psi_{\mathrm{c,i}}\in\{$$,$$\}$ 33.3  ° , 50.26  ° }, where the lower bound is selected from the $3\sigma$ confidence interval from the observational posteriors, and the upper bound from the $2\sigma$ confidence interval of our four-body scattering simulations. The asymmetric lower bound is selected to encompass more of the highly non-Gaussian observational posterior. The nodes are randomly distributed. HAT-P-11 b is initialized with orbital elements randomly drawn from uniform distributions informed from the $2\sigma$ distributions obtained in our four-body scattering simulations: $a_{\mathrm{b}}=\{0.154\ \mathrm{AU},0.518\ \mathrm{AU}\}$, $e_{\mathrm{b}}=\{0.01,0.70\}$ and $\psi_{\mathrm{b}}=\{1.01^{\circ},97.0^{\circ}\}$. We initialize the mass with the best-fit value from Table 1, and the radius with Equation (13).
We note that our scattering simulations often include an extra surviving distant planet. The impact of distant perturbers on the classic ZLK hierarchical configuration was studied by Best & Petrovich (2022), who found that it generates chaotic spin-orbit evolution and makes achieving retrograde orbits easier, which would be favorable to our scenario. However, the first-order dynamics do not change, so we do not model it in our simulations. We also comment on the possibility of direct ZLK oscillations occuring between HAT-P-11 c and this surviving distant perturber. This is a fairly rare occurrence; only 653 simulations, or 14%, result in a mutual inclination between HAT-P-11 c and the surviving outer planet greater than 40°40°40  °, so ZLK oscillations are rarely triggered. However, for these simulations where the prerequisite mutual inclination for ZLK cycles is reached, $|\dot{\omega}_{\mathrm{ZLK}}|>|\dot{\omega}_{\mathrm{GR}}|$ so ZLK cycles are in fact active. Nevertheless, the parameter space for a surviving external planet is large and unconstrained from observation, and it doesn’t change the dynamics of the inner planets qualitatively. Thus, we do not model this outer planet in our ZLK simulations.
@@
We also model the effects of general relativity and tides. We use REBOUNDx (Tamayo et al., 2020) to model both. For GR precession, we use the gr prescription (Anderson et al., 1975), appropriate for systems with a dominant central mass. We consider self-consistent spin, tidal and orbital evolution using the tides_spin effect (Lu et al., 2023), which provides the ability to evolve the evolution of the stellar spin axis for consistent spin-orbit misalignment tracking as well. For planet b, we use fiducial parameters of $k_{2}=0.5$, $I=0.25$ appropriate for a Uranus-like planet (Yoder, 1995; Lainey, 2016), and we initialize its spin state as both pseudo-synchronized and aligned with the orbit. Planet c is considered a point particle, and the effects of structure are not accounted for. While evolution of the stellar spin axis is considered as part of the equilibrium tides model, spin-down due to stellar evolution, which would occur during the early phases of our simulations (Bouvier et al., 1997), is not modeled. These effects are not expected to be significant for our simulations. Authors such as Bolmont et al. (2012) and Faridani et al. (2023, 2024) explored the effects of stellar spin evolution on planetary orbits and only found meaningful impacts for planets within $a=0.05$ AU, well within the hypothesized original orbit of HAT-P-11 b.
@@
Tides on the host star and planet c are not considered, but we set the value of planet b’s tidal parameter to $\tau_{\rm b}=10^{-5}$ years. This is an unrealistically high value of $\tau$, chosen to minimize computation time: direct N-body simulations with a realistic $\tau$ value proved to be computationally infeasible. This is a common practice in numerical simulations (e.g Bolmont et al., 2015; Becker et al., 2020). Scaling $\tau$ does not change the qualitative behavior of the system. Neither the ZLK timescale nor the tidal precession rate, Equations (6) and (12) respectively, depend on the value of the tidal parameter. The only dependence on $\tau$ is the rate of orbital energy drained from tides. Thus, our scaled-up value of $\tau$ serves only to drain more energy from the planet’s orbit every ZLK cycle (therefore decreasing the semimajor axis more) and increase the rate at which HAT-P-11 b circularizes once the ZLK oscillations are ultimately quenched. These both only serve to linearly modify the timescale in which the planet reaches the final stage of evolution.
@@
We now briefly discuss what realistic values of $\tau$ may be. While the values of tidal parameters in exoplanets are very poorly constrained, estimates of Neptune’s $\tau\sim 10^{-8}$ years are based on the orbit of its satellites Proteus and Larissa (Zhang & Hamilton, 2007). Therefore, the tidal timescale in the N-body simulations is a factor $\sim$103 longer than the simulation timescale.
@@ 3.5 Simulation Results
Figure 7: Results from a population synthesis of ZLK simulations. We consider $40$ simulations with radius evolution (orange) and $40$ with fixed radius (blue). We compare distributions of semimajor axis, spin-orbit misalignment and mutual inclination. We see that the fixed radius model cannot generate HAT-P-11 b’s present-day semi-major axis, but the model with tidally-driven radius inflation can.
@@
Figure 7 shows distributions of spin-orbit misalignment and eccentricity for a population synthesis of $100$ systems. We initialize each simulation as described in the previous section and run until the ZLK cycles are quenched and the orbit has circularized to the present-day eccentricity $e_{\mathrm{b}}=0.218$. At this point, the simulation is halted. To illustrate the effects of radius inflation, we also run the same initial condition but fix the radius of HAT-P-11 b to its present-day value throughout the course of the simulation. We allow each simulation to run for 10 Myr, which corresponds to 10 Gyr in when time is rescaled appropriately. $40$ simulations with radius inflation are able to quench ZLK oscillations and circularize in this timeframe, with the remaining $60$ either never initiating ZLK oscillations or being unable to quench them in time. For the simulations with fixed radius, $40$ simulations quench and circularize in time. For the population synthesis that considers radius inflation, we see that while the best-fit observational values are not the most favored there is nonetheless a high proportion of simulations that match the constraints, and we do not require fine-tuning to reproduce our posteriors. The population synthesis without radius inflation does not exhibit significant differences in the spin-orbit misalignment or mutual inclination distributions. However, as expected this model greatly underpredicts the semimajor axis of HAT-P-11 b: all simulations predict an orbit with $a_{\mathrm{b}}<0.045$ AU (consistent with the implication of Figure 5).
@@
Figure 8: Fiducial ZLK migration of HAT-P-11 b. From top to bottom, HAT-P-11 b’s semimajor axis, eccentricity, spin-orbit misalignment and radius are shown with the colored lines. The black dots represent HAT-P-11 b’s measured orbital parameters, with the error bar indicating the $1\sigma$ confidence interval. We posit that HAT-P-11 b is still in the tidal circularization stage after the suppression of ZLK oscillations. There is excellent agreement between our simulation results and observed values.
@@
Figure 8 shows one of our simulations in detail. Table 2 lists the initial orbital parameters of the simulation, as well as the final values of of interest. During the first phase of the simulation, ZLK oscillations are active and the planet experiences the characteristic coupled eccentricity and inclination oscillations. At each high-eccentricity epoch, tides act on the planet to lower its semimajor axis. The radius evolution of the planet is also plotted, and we see that during each high-eccentricity epoch the planet inflates to over twice its present-day radius. Note that the minimum pericenter distance (${\sim}0.02\ \mathrm{AU}$) is well outside twice of the tidal disruption radius ($r_{t}=R_{p}(M_{*}/M_{p})^{1/3}{\sim}0.008\ \mathrm{AU}$), so we do not include mass loss due to tidal stripping in our model, as the effects are weak (Guillochon et al., 2011). Later, the ZLK oscillations are damped during a high eccentricity epoch. Tides continue to act to circularize the planet’s orbit. In the meantime, the orbit normal of HAT-P-11 b precesses about the invariant plane of the system (essentially HAT-P-11 c’s orbit normal), as expected, resulting in periodic spin-orbit misalignment oscillations. As expected, the stellar spin axis and the orbit of HAT-P-11 c do not change significantly over the course of the integration.
@@
We comment again on the interpretation of timescales, given our choice of $\tau$ in order to speed up computations. Please refer to Section 3.4 for justification regarding the choice of an enhanced dissipation rate. The enhanced dissipation rate affects the phase when tidal dissipation dominates, and thus we scale the timescales on the right-hand side of the panel to the realistic values accordingly. Note that with a more realistic tidal time-lag, we would expect to see more spin-orbit misalignment cycles on the right-hand side of the plot when tides dominate (since the orbital precession frequency does not depend on tidal dissipation and shouldn’t be scaled). In addition, we would expect to have more ZLK cycles on the left-hand side, since each cycle leads to a weaker semi-major axis decay. However, the qualitative behavior of the evolution is the same.
@@
The age of the HAT-P-11 system is not well-constrained: Yee et al. (2018) propose an age of $6.5^{+5.9}_{-4.1}$ Gyr, while Morton et al. (2016) present an estimate of $2.69^{+2.88}_{-1.24}$ Gyr. Our results are consistent with both estimates of the system age. Note that the timescale of ZLK migration varied by around an order of magnitude, depending on initial conditions.
@@
$\mathrm{R}$ ($\mathrm{R_{Earth}}$) 3.69 4.27 N/A N/A
$a$ (AU) 0.19 0.056 4.1 4.1
$e$ 0.48 0.22 0.65 0.65
$\psi$ $(^{\circ})$ 82 156 35.3 35.4
Note. — Relevant input parameters and output values of the REBOUND simulation shown in Figure 8. The final values are the values at the time corresponding to the black dot in Figure 8, which is where $e_{\mathrm{b}}$ is equal to the best fit value. The orbit of HAT-P-11 c does not change significantly over the course of the simulation.
The present-day observational constraints on HAT-P-11 b’s orbit are also shown on Figure 8. Our simulation is able to match all values very well: we reproduce all observed values within $2\sigma$ other than semimajor axis, which is well-reproduced qualitatively. We propose that HAT-P-11 b is currently in this final stage of evolution and is still in the process of tidal circularization. We conclude that ZLK migration coupled with thermally driven radius evolution is capable of explaining the present-day orbit of HAT-P-11 b remarkably well. Again, we emphasize that we did not perform an exhaustive parameter space analysis of this system. Rather, our aim is to qualitatively show that such a process is capable of reproducing HAT-P-11 b’s orbit.
[…]